OpenStudy (anonymous):

he sum of three numbers is 26. The second number is twice the first and the third number is 6 more than the second. Find the numbers.

OpenStudy (anonymous):

Let the first no. be x. then 2nd no. = 2x and 3rd no. = 2x+6 Now, x+(2x)+(2x+6)=26 5x=26-6 5x=20 x=4 first no. = 4 2nd no. = 2x = 8 3rd no. = 2x+6 = 14

OpenStudy (anonymous):

Let the second number be: s Let the first number be : f Le the third number be : t So, s = 2f t = 6 + s So, f + s + t = 26 s/2 + s + 6 + s = 26 5s + 12 = 52 Find s from here..